The Wood Pile Due on Thursday, Feb 7 at 11:59 pm (EST) (clp18) A cord is a volume of cut wood equal to a stack 4ft high, ) equals one cord? (Note: When entering the answer with a unit that has a power such as this one, you must type m circumflex 3. Circumflex is shift-6 on your keyboard.) 4ft wide, and 8ft long. How many cubic meters (m3 Tries 0/8
Answer #1

As 1 foot = 0.305 meters

Hence 4 ft = 1.22 m and 8 ft = 2.44 m

The cord is just like cuboid of length (l) = 2.44 m ,width (w)=1.22 m and height (h)=1.22 m

volume of cuboid = l*b*h

so volume of code = 2.44*1.22*1.22  = 3.631 m^3

hence one cord = 3.631 m^3